# Beyond the Silence: How Men Navigate Infertility Through Digital Communities and Data Sharing

### v1

@@ 4.1.3. Moderation: Structure, Orientation, and Safety
4.1.4. Predictors of More Future Engagement
The time-lagged regression analysis presented in Tables 1-3 offers insights into how language use (sentiment), themes discussed by the user, and user engagement behavior on this subreddit community predict higher engagement from month to month in the intfertiility support community. Specifically, linguistic markers drawn from LIWC (Table 1) show that future engagement and language showing affiliation with others in the community, insight-oriented words, and cognitive processing terms are all positively associated with an increased engagement in the community. Conversely, language associated with detachment or externalization—such as references to money, negation, or fatalistic expressions—is negatively associated with continued participation.
@@ 4.1.4. Predictors of More Future Engagement
4.2. Community Structure and User Role Dynamics
@@
Indeed, we struggled to find thematic signatures for different clusters throughout our qualitative analysis of the clusters. With the exception of smaller clusters like cluster 2 made of 33 users, which focused on at-home sperm banking, most clusters showed a similar spread of varied topics related to male infertility. Even in that cluster, the discussion diverged to multiple different topics. However, similarities and differences emerged between different roles, and the ways in which they were enacted in different clusters. We discuss this in detail in §4.2.1.
@@ 18 Mentor; Experts; Systems Critic; Hope Broker
4.2.1. Cross-Cluster Roles and Their Variation
Certain roles recur across clusters but manifest differently depending on the context. These roles are overlapping – often one poster is wearing “many hats” frequently within the context of a given thread as they balance providing medical information with socializing markers of commiseration. Users in these roles adopt certain commenting behaviors depending upon which role they are undertaking. Table 5 shows the breakdown of roles in different clusters.
Mentors, who represent the majority of the dominant voices within our clusters, are those participants who are ”shepherding” new and old members alike. Much of this guidance comes in the form of emotional support and advice based on their own lived experience. In Cluster 7, one mentor described in detail expected recovering sperm count thresholds after hormonal therapy, while in Cluster 6 another narrated their journey with acceptance that they would not have a child after multiple failed surgeries. In both examples, these Mentors modeled to other participants how they faced similar struggles with fertility: the former through detailed experiential medical information about recovery, and the latter through closure after failure. Mentors acting as hope brokers are nearly universal, sustaining morale across contexts: either by framing setbacks in terms of relatable shared experiences or by providing medical guidance to overcome obstacles. In Cluster 18, one partner emphasized improvements after months of supplements, while in Cluster 23 another highlighted dramatic gains following a vitamin regimen. Such accounts sustain optimism even when treatment outcomes remain uncertain.
Experts in different clusters offered detailed instructions on hormone regimens, supplements, and surgical timing. In Cluster 5, a user explained how Clomid influences hormonal pathways. In Cluster 11, another laid out strategies for recovering after testosterone therapy; and in Cluster 15, individuals logged precise hormone dosages and outcomes. Across clusters, experts provide structured, technical guidance using sometimes mechanistic or even clinical language, sprinkled occasionally with their anecdotal observations. Having extensive knowledge of an often opaque and byzantine landscape of navigating infertility health challenges, these users highlight the challenges all these users are facing. As such, comments that reflect their frustration with these health systems punctuate nearly every cluster. This systems critical behavior is often reflected by both the Expert and Mentor roles, as their first hand experience of the structural bias, financial burden, and professional blind spots makes their insights carry more weight. For instance, one user argued that fertility clinics often marginalize men, another criticized doctors for pressuring patients into surgery, and others warned about conflicts of interest in treatment recommendations. System critical comments thus shape the community’s positioning against the medical establishment, making visible the broader structures that frame individual infertility struggles. These communities exist largely because of the hardships their members have faced accessing and affording care as well as overcoming structural and societal biases regarding male infertility; it is no surprise that an undercurrent of systems criticism is so clearly visible.
@@ 4.2.2. Moderation by Users Across Clusters
Human moderators and everyday users enforce and interpret group norms via often explicit forum guidelines for post behavior. Some users apply the rules strictly, telling others that their medical results do not qualify for posting. Others soften the enforcement, explaining why certain parameters (such as morphology alone) are not considered meaningful when posting their medical records and asking for insights. Still others frame moderation through community response, pointing out how upvotes and engagement reflect shared perceptions of value of certain posts – e.g., posts by system critical users critiquing the medical establishment through the lens of their experiences. Together, these practices show how moderation is enacted both through both automation and social interpretation.
In Cluster 4, we identified an instance of norm-setting through empathy, as users caution against flaunting “good” results that might trigger others in the community. We found multiple examples of moderation taking the form of scientific correction, with members pushing back against poor medical advice and urging others to seek proper specialists. However, we found that Cluster 23 had little formal moderation; extreme claims, such as abstinence-based fertility hacks were not checked. Cluster 23 is one of the smallest identified clusters with 26 total users. This indicates that smaller clusters might have different moderation styles than the rest of the community.
Moderation thus emerges as a form of community guidance, balancing rule enforcement with “navigation labor.” This labor, often invisible, ensures that newcomers find their footing and that discussions remain accessible, high-signal, and emotionally safe. Importantly, mentors sometimes overlap with moderators in this function, herding discussions in ways that preserve both coherence and empathy.
Figure 2 summarizes the different roles we identified in our analysis. It shows that mentors and moderators do not only enforce posting standards, but actively perform way-finding by directing users to appropriate threads, providing structured resources, and sometimes intervening with empathetic language to soften redirects. By doing so, they serve as way-finders, moving discussions to appropriate threads, orienting newcomers, and preventing newcomers from feeling “lost” when first engaging in the subreddit. This shepherding role is not purely administrative—it blends procedural guidance with emotional counseling, often drawing on lived experiences to normalize confusion. For example, mentors not only provide medical or experiential knowledge but also explicitly guide members toward appropriate threads, FAQs, or resources, ensuring that vulnerable disclosures are not left unanswered. This highlights the dual role of mentors and moderators as both emotional anchors and navigators of community norms.
@@ 4.2.3. The Role of Women and Significant Others Across Clusters
Women and partners are especially visible in clusters where emotional processing and distrust of the medical system dominate. In Cluster 4, women acted as voices of empathy and shared grief. For example, one partner expressed sorrow over a recent miscarriage and directed others toward supportive communities, simultaneously affirming shared pain and providing peer guidance.
In Cluster 18, wives frequently narrate their husbands’ test results and frustrations with medical encounters. One described her partner’s severe semen analysis results, while another, in her capacity as a system critic recounted a doctor dismissing a possible connection between varicoceles (a medical condition that causes pain and swelling in the testicles) and infertility.
In other clusters, partners take on the expert role by being “partner-scientists.” They compile supplement regimens and interpret semen analysis results on behalf of their husbands. For example, one partner described how her husband’s sperm count continued to rise as he adhered to lifestyle changes, while another detailed an exact supplement stack that reportedly improved results. These examples illustrate how women take on the roles of record-keepers, advocates, and interpreters of their partners’ medical journeys in different clusters.
@@
4.3. Cross-Posters as Navigators and Mentors in the Male Infertility Community
@@
Cross-posters on Reddit can be characterized by both their activity patterns and linguistic style. They are generally highly active users, contributing more, longer comments than non-cross-posting users, with shorter average response times, and longer, more positive responses, indicating acceptace by other members of the community.
Comments by cross-posters usually focus on navigation, where mentors and moderators (see §4.2.2) help users to navigate (see §4.1.3 the norms of r/maleinfertility. Some of the moderators rely on their knowledge of the normative differences between r/maleinfertility and other subreddits to guide users through its use. In addition to guiding people in community navigation, cross-posting users also guided others in better understanding their medical reports (see §4.1.1).
@@ 5.1. Contributions to Current Literature
We organize our contributions around three analytic dimensions. First, affiliation, insight, and actionable guidance drive sustained engagement through informal diagnostics, logistical coordination, and therapeutic scaffolding. Second, cohesive micro-communities with adaptive, overlapping roles show how mentorship and moderation extend peer-support and platform-governance accounts. Third, cross-posters normalize norms and transfer know-how across communities, complicating models of knowledge brokering and collective resilience. Together these findings advance theories of social exchange, trauma-informed design, and platform governance by foregrounding the hybrid infrastructures and relational dynamics that sustain participation in stigmatized health contexts.
@@ 5.1.1. Affiliation, insight, and actionable guidance keep people coming back
To answer RQ1, we combined thematic analysis (115 topics; §4.1) and time-lagged regression (§4.1.4). Three discourse arenas emerged—Medical Advice (63.3%), Emotional Support (7.4%), and Moderation/Orientation (29.2%)—forming a hybrid infrastructure for diagnostics, therapy, and governance. Engagement rose with actionable guidance but declined when users received more replies, countering earlier findings that community responses foster future participation (Arguello et al., 2006; Lampe and Johnston, 2005; Backstrom et al., 2008; Chung et al., 2017; Lampe et al., 2010; Hwang and Foote, 2021). We unpack these patterns below.
Informal diagnostics and logistics.
@@
The dominance of Medical Advice shows how peers translate lab values, treatment options, and system navigation into lay expertise that supplements—or compensates for—opaque clinical interactions. This extends work on socially sustained self-tracking and collective sensemaking, where personal data becomes shared, emotionally manageable, and actionable narratives (Li et al., 2010; Knittel et al., 2021; Mamykina et al., 2015; Choe et al., 2017; Lu et al., 2021; Chung et al., 2017; Gui et al., 2017). Social Exchange Theory explains why such disclosures persist despite stigma: intrinsic rewards (self-worth, reciprocity) and informational gains outweigh perceived risks (Emerson, 1972; Wu et al., 2023; Yan et al., 2016; Soc, 2011).
Therapeutic scaffolding
Emotional Support. Emotional Support spans grief processing and “comfort posting,” aligning with research on digital belonging and men’s mental health, where social connection mitigates depression and suicidality (McLaren and Challis, 2009; Bailey et al., 2018; Holt-Lunstad and Steptoe, 2022; Steptoe et al., 2013). Asynchronous communication can likewise foster well-being (Hull et al., 2020; Rahman and Nayan, 2022) and support resilience and identity repair in trauma-adjacent settings (Ammari et al., 2025; Randazzo and Ammari, 2025). Yet our regression analysis complicates the common assumption that more replies drive retention: response probability negatively predicts subsequent activity. Possible explanations include overly corrective or impersonal feedback and off-platform migration (e.g., DMs, external groups), revising participation models in stigmatized contexts (Andalibi, 2020; Ammari et al., 2019).
@@
While prior work highlights direct peer responses as central to support (Hwang and Foote, 2021)—a key tenet of trauma-informed design (Chen et al., 2022; Scott et al., 2023)—users may instead sustain engagement by lurking (Randazzo and Ammari, 2023) or distributing participation across multiple communities (Hwang and Foote, 2021; Ammari et al., 2025). We outline implications for these dynamics in the future work section.
@@ 5.1.2. Cohesive micro-communities with adaptive and overlapping roles
Community detection revealed structurally cohesive clusters that fell well below topical/stance purity thresholds, indicating they are not echo chambers (Morini et al., 2021). Within these clusters, shared and distinct roles emerged, expressed differently across groups.
@@
r/maleinfertility is tightly moderated, as evidenced by its About section on the homepage which directed partners (of infertile men) to a dedicated thread. Moderation—both automated and human—guides information flow by redirecting posts to wikis or daily threads and safeguards the psychosocial well-being of community members by setting rules about acceptable content. Yet these same controls can alienate newcomers, and it remains unclear which moderation practices dampen engagement. This reflects the contextual nature of harm online (Gillespie, 2018): even well-intentioned “good news” posts may be discouraged when negative sentiment sustains participation (§4.1.4). This dynamic parallels the “pain olympics” among parents of children with special needs, where positive updates can provoke negative responses from peers facing deeper challenges (Ammari and Schoenebeck, 2015).
@@
While peer support is a recognized cornerstone of trauma-informed design (Chen et al., 2022; Scott et al., 2023), our findings show that support extends beyond upvotes, replies, or algorithmic curation (LaViolette and Hogan, 2019; Randazzo and Ammari, 2023). Mentors and moderators play critical way-finding roles: contextualizing rules, directing newcomers to curated resources, and reframing emotional disclosures into actionable steps. Building on platform governance and care-work scholarship, we identify an under-theorized “navigation labor,” where these actors blend rule enforcement with empathetic guidance.
@@
Other roles that overlap with moderation—particularly mentorship—complicate conventional models of peer support. The type of peer who engages with newcomers and how these interactions evolve shape both emotional safety and knowledge transfer, suggesting a fluid, multi-layered ecology of support.

### v2

@@ 4.2. Predictors of More Future Engagement (RQ1b)
4.3. Community Structure (RQ2a)
@@
Indeed, we struggled to find thematic signatures for different clusters throughout our qualitative analysis of the clusters. With the exception of smaller clusters like cluster 2 made of 33 users, which focused on at-home sperm banking, most clusters showed a similar spread of varied topics related to male infertility. Even in that cluster, the discussion diverged to multiple different topics. However, similarities and differences emerged between different roles, and the ways in which they were enacted in different clusters. We discuss this in detail in §4.4.
@@
4.4. Cross-Cluster Roles and Their Variation (RQ2b)
Certain roles recur across clusters but manifest differently by context. These roles overlap—one poster often wears many hats within a thread, balancing medical information with markers of commiseration. Table 5 shows role distribution across clusters. Mentors, representing the majority of dominant voices, shepherd new and established members through emotional support and lived-experience advice. In Cluster 7, one mentor detailed expected sperm count recovery thresholds after hormonal therapy; in Cluster 6, another narrated their journey toward acceptance after multiple failed surgeries. Both modeled how to face fertility struggles—the former through experiential medical information, the latter through closure after failure. Mentors acting as hope brokers are nearly universal, sustaining morale by framing setbacks as relatable shared experiences or providing guidance to overcome obstacles. In Clusters 18 and 23, partners emphasized improvements following supplement and vitamin regimens, sustaining optimism even when outcomes remained uncertain. Experts offered detailed instructions on hormone regimens, supplements, and surgical timing. In Cluster 5, a user explained how Clomid influences hormonal pathways; in Cluster 11, another outlined testosterone therapy recovery strategies; in Cluster 15, individuals logged precise dosages and outcomes. Experts provide structured, technical guidance using clinical language, occasionally sprinkled with anecdotal observations. Both Expert and Mentor roles frequently overlap with system critical behavior, as firsthand experience of structural bias, financial burden, and professional blind spots lends weight to their insights. Users argued that fertility clinics marginalize men, criticized doctors pressuring patients into surgery, and warned about conflicts of interest. These communities exist largely because members have faced hardships accessing and affording care while overcoming societal biases regarding male infertility—an undercurrent of systems criticism is unsurprising.
@@ 4.4.1. Moderation by Users Across Clusters
Human moderators and everyday users enforce group norms via explicit forum guidelines. Some apply rules strictly, telling others their results do not qualify for posting. Others soften enforcement, explaining why certain parameters (such as morphology alone) are not meaningful. Still others frame moderation through community response, noting how upvotes reflect shared perceptions of value. Together, these practices show moderation enacted through both automation and social interpretation.
In Cluster 4, we identified norm-setting through empathy, as users cautioned against flaunting good results that might trigger others. Multiple examples showed moderation as scientific correction, with members pushing back against poor advice and urging others to seek specialists. However, Cluster 23—one of the smallest with 26 users—had little formal moderation; extreme claims like abstinence-based fertility hacks went unchecked, indicating smaller clusters may have different moderation styles. Moderation thus emerges as community guidance balancing rule enforcement with navigation labor—often invisible work ensuring newcomers find their footing and discussions remain accessible, high-signal, and emotionally safe. Mentors sometimes overlap with moderators, herding discussions to preserve both coherence and empathy.
Figure 2 summarizes identified roles. Mentors and moderators not only enforce posting standards but perform way-finding: directing users to appropriate threads, providing structured resources, and intervening with empathetic language to soften redirects. This shepherding blends procedural guidance with emotional counseling, drawing on lived experience to normalize confusion and ensuring vulnerable disclosures are not left unanswered.
@@ 4.4.2. The Role of Women and Significant Others Across Clusters
Women and partners are especially visible in clusters where emotional processing and medical distrust dominate. In Cluster 4, women acted as voices of empathy and shared grief—one partner expressed sorrow over a miscarriage and directed others toward supportive communities, affirming shared pain while providing peer guidance.
In Cluster 18, wives frequently narrated their husbands’ test results and frustrations with medical encounters. One described her partner’s severe semen analysis results; another, as system critic, recounted a doctor dismissing a possible connection between varicoceles [^7] A varicocele is an enlargement of the veins in the scrotum that drain the testicles, which can lead to pooling of blood and may negatively affect male fertility by increasing testicular temperature and impairing sperm production – Source: https://www.mayoclinic.org/diseases-conditions/varicocele/symptoms-causes/syc-20378771 and infertility. In other clusters, partners became ”partner-scientists,” compiling supplement regimens and interpreting semen analyses on their husbands’ behalf. One described how her husband’s sperm count rose with lifestyle changes; another detailed an exact supplement stack reportedly improving results. These examples illustrate women’s roles as record-keepers, advocates, and interpreters of their partners’ medical journeys—turning intimate partnerships into sites of experimentation balancing hope with vigilance.
@@
4.5. Cross-Posters as Navigators and Mentors in the Male Infertility Community (RQ3)
@@
Cross-posters on Reddit can be characterized by both their activity patterns and linguistic style. They are generally highly active users, contributing more, longer comments than non-cross-posting users, with shorter average response times, and longer, more positive responses, indicating acceptance by other members of the community.
Comments by cross-posters usually focus on navigation, where mentors and moderators (see §4.4.1) help users to navigate (see §4.1.3) the norms of r/maleinfertility. Some of the moderators rely on their knowledge of the normative differences between r/maleinfertility and other subreddits to guide users through its use. In addition to guiding people in community navigation, cross-posting users also guided others in better understanding their medical reports (see §4.1.1).
@@ 5.1. Contributions to Current Literature
We organize our contributions around three analytic dimensions. First, reflecting on RQ1, §5.1.1 explores how affiliation, insight, and actionable guidance drive sustained engagement through informal diagnostics, logistical coordination, and therapeutic scaffolding. Second, by reflecting on the findings associated with RQ2, §5.1.2 explore cohesive micro-communities with adaptive, overlapping roles and show how mentorship and moderation extend peer-support and platform-governance accounts. Third, reflecting on RQ3, §5.1.3 explores how cross-posters normalize norms and transfer know-how across communities, complicating models of knowledge brokering and collective resilience. Together these findings advance theories of social exchange, trauma-informed design, and platform governance by foregrounding the hybrid infrastructures and relational dynamics that sustain participation in stigmatized health contexts.
@@ 5.1.1. Affiliation, insight, and actionable guidance keep people coming back
Based on thematic analysis (115 topics; §4.1) and time-lagged regression (§4.2), three discourse arenas emerged—Medical Advice, Emotional Support, and Moderation/Orientation —forming hybrid infrastructure for diagnostics, therapy, and governance. Engagement increased with actionable guidance but declined when users received more replies, countering findings that community responses foster future participation (Arguello et al., 2006; Lampe and Johnston, 2005; Backstrom et al., 2008; Chung et al., 2017; Lampe et al., 2010; Hwang and Foote, 2021).
Informal diagnostics and logistics: Challenging Traditional Masculinity Ideologies
@@
The dominance of Medical Advice shows how peers translate lab values, treatment options, and system navigation into lay expertise supplementing opaque clinical interactions. This extends work on socially sustained self-tracking and collective sensemaking, where personal data becomes shared, emotionally manageable narratives (Li et al., 2010; Knittel et al., 2021; Mamykina et al., 2015; Choe et al., 2017; Lu et al., 2021; Chung et al., 2017; Gui et al., 2017). Social Exchange Theory explains why disclosures persist despite stigma: intrinsic rewards and informational gains outweigh perceived risks (Emerson, 1972; Wu et al., 2023; Yan et al., 2016; Soc, 2011).
Yet technical framing obscures the emotional substrate. Viewing digital practices through a sociomaterial lens—where identities are enacted and negotiated across spaces through material and discursive means (Van House, 2011; Barad, 2007)—reveals how members simultaneously challenge and comply with Traditional Masculinity Ideologies (TMI). Much as DIY fathers appropriate traditionally feminine caregiving by reframing it through masculine logics of problem-solving and technical competence (Ammari et al., 2017), men in r/maleinfertility embed emotional support within medical discourse: decoding semen analyses, comparing protocols, and crowdsourcing interpretations. This allows vulnerability to circulate under the guise of information exchange—legitimizing grief often overlooked offline without explicitly violating TMI norms demanding stoicism (Messerschmidt, 2019; Cleary, 2012).
This explains why members characterized r/infertility as ”female-centered” and ”emotionally focused” while viewing r/maleinfertility as data-driven—despite both providing substantial emotional support. The difference lies not in care work’s presence but its discursive packaging: r/maleinfertility members performed support through technical registers preserving masculine self-presentation. Dolan et al. (Dolan et al., 2017) observed similar patterns clinically, where men defined themselves as the ”voice of reason” and characterized emotional expression as something women did—limiting disclosure while positioning stoicism as masculine rationality. The medical framing served an analogous function, enabling men to ”work around” the ”perverse passivity” Dolan identified—where infertility was stoically endured rather than actively addressed—by channeling emotional processing through diagnostic sensemaking.
While this technical framing mirrors women’s infertility communities in emphasizing information support (Zou et al., 2024)2.2.2), r/maleinfertility demonstrated robust esteem support, particularly around identity-threatening experiences like donor conception. The insistence that men using donors are not ”lesser men because of a medical issue”4.1.2) directly addresses the existential crisis Dolan (Dolan et al., 2017) documented, where infertility ”destabilised the ontological connections between a prized masculine identity, the male body and biological fatherhood.” Where women’s communities offered informational solidarity while avoiding deep relational investment, the male community provided explicit identity repair—reconstructing fatherhood around influence and presence rather than biology.
This resonates with Semaan et al.’s findings on military veterans, where hyper-masculinity learned during service created barriers to post-service disclosure (Semaan et al., 2016, 2017). Veterans who successfully navigated transition often did so by witnessing others disclose online, establishing ”a new form of camaraderie not based solely on a hyper-masculine identity” (Semaan et al., 2017). Similarly, r/maleinfertility members observing peers share vulnerable experiences within technical frameworks appeared more willing to disclose—suggesting the medical register provided a TMI-compliant pathway to emotional expression. Hegemonic masculinity operates through cultural ascendancy rather than overt domination (Messerschmidt, 2019); the technical discourse may represent a ”complicit masculinity” realizing some benefits of masculine norms while enabling care work and esteem support that sustains the community and facilitates identity repair.
Therapeutic scaffolding.
Emotional Support spans grief processing and ”comfort posting,” aligning with research on digital belonging and men’s mental health, where social connection mitigates depression and suicidality (McLaren and Challis, 2009; Bailey et al., 2018; Holt-Lunstad and Steptoe, 2022; Steptoe et al., 2013).
@@
Additionally, asynchronous communication fosters well-being (Hull et al., 2020; Rahman and Nayan, 2022) and supports resilience in trauma-adjacent settings (Ammari et al., 2025; Randazzo and Ammari, 2025). Yet our regression complicates the assumption that more replies drive retention: response probability negatively predicts subsequent activity. Possible explanations include overly corrective feedback and off-platform migration, revising participation models in stigmatized contexts (Andalibi, 2020; Ammari et al., 2019). While prior work highlights direct peer responses as central to support (Hwang and Foote, 2021)—a tenet of trauma-informed design (Chen et al., 2022; Scott et al., 2023)—users may sustain engagement through lurking (Randazzo and Ammari, 2023) or distributing participation across communities (Hwang and Foote, 2021; Ammari et al., 2025). We outline implications in the future work section.
@@ 5.1.2. Cohesive micro-communities with adaptive and overlapping roles
Community detection revealed structurally cohesive clusters below topical purity thresholds—not echo chambers (Morini et al., 2021)—yet the network exhibited notable structural closure (§4.3) with substantial reciprocity, while maintaining topical breadth. This complicates claims that pseudonymous support fosters weak, transactional ties (Zou et al., 2024), and shows that they can provide network support (see §2.2.2). However, such closure may paradoxically impede healing given how tight relational networks can produce grief bubbles that trap users in trauma-focused interactions, hindering identity integration even when content remains diverse (Randazzo and Ammari, 2025). The scarcity of male-centric infertility spaces (Patel et al., 2019) may intensify this by concentrating participation into relationally dense communities (we discuss a temporal element to this issue in future work). We next discuss the roles we discovered in this community.
@@
Wu et al. (Wu et al., 2024) show how moderators establish symbolic boundaries differentiating in-group from out-group members. In r/maleinfertility, this manifested in discussions defining who qualifies as clinically infertile versus facing general conception challenges, with the subreddit’s About section directing partners to dedicated threads. As detailed in §4.1.3, moderators specified that men with legitimate infertility claims (as opposed to low libido) qualify for membership, though some clusters proved more permissive than others (4.3). Unlike fathers-only groups that explicitly excluded mothers (Ammari and Schoenebeck, 2016), r/maleinfertility’s exclusion criteria remained ambiguous. This nuanced boundary-setting at the intersection of gender identities (Gilbert, 2023) and health conditions (Saha et al., 2020) introduced friction that complicated safe space management. For instance, in Cluster 4, users cautioned against flaunting good results that might trigger others—paralleling ”pain olympics” among parents of children with special needs (Ammari and Schoenebeck, 2015). This reflects the contextual nature of online harm (Gillespie, 2018): even well-intentioned ”good news” posts may be discouraged when negative sentiment sustains participation (§4.2).
@@
Peer support is central to trauma-informed design (Chen et al., 2022; Scott et al., 2023), and our findings show support extends beyond upvotes and replies (LaViolette and Hogan, 2019; Randazzo and Ammari, 2023). Mentors and moderators play critical way-finding roles: contextualizing rules, directing newcomers to resources, and reframing emotional disclosures into actionable steps. This information management—redirecting posts to wikis, daily threads, or pinned resources—fulfills a key moderation function (organizing information (Edwards, 2002)) but can frustrate newcomers unfamiliar with community architecture (§4.1.3). Gilbert (Gilbert, 2023) demonstrates how such moderation constitutes intensive emotional labor at personal, community, and systemic levels. Building on this, we identify an under-theorized ”navigation labor” where actors blend rule enforcement with empathetic guidance, shepherding vulnerable disclosures so they are not left unanswered.
@@
Roles overlapping with moderation—particularly mentorship—complicate conventional peer support models and reveal moderation as simultaneously protection and collective empowerment (Ammari et al., 2022). Two roles warrant attention: system critics and hope brokers. System critics—users critiquing fertility clinics for marginalizing men and warning about conflicts of interest—perform what Ammari et al. (Ammari et al., 2022) term ”empowerment as consciousness” (power within): questioning institutional structures upon finding others facing similar constraints. This criticism reflects material reality: research consistently excludes men from fertility studies despite fertility being couple-based (Harlow et al., 2020), leaving users ”grasping at straws” for male-factor information (§4.1.1).
Hope brokers complement this by enacting ”empowerment through community” (power through)
—members gaining resources and efficacy through experienced peers (Ammari et al., 2022). By framing setbacks as relatable and highlighting treatment gains, they sustain optimism when outcomes remain uncertain, mirroring ”networked empowerment” among parents of children with special needs (Ammari and Schoenebeck, 2015). Together, system critics raising consciousness and hope brokers modeling pathways create conditions for ”empowerment in community” (power with): collective action addressing common concerns (Ammari et al., 2022). The overlapping roles (Figure 2) reveal users wearing ”many hats” within single threads—distributed governance building collective capacity to challenge inadequate care systems and hegemonic masculinity (Messerschmidt, 2019; Dolan et al., 2017). Understanding peer support evolution is central to trauma-informed design (§2.3).